alishia:

A new coffee shop can hold no more than 50 seats. The owner wants at least 20 of the seats to be stools and the remaining seats to be recliners. If x is the number of stools and y is the number of recliners, which graph represents the solution to the system of inequalities?

alishia:

please help me i am dying because of this question

Nnesha:

there are 2 situations: 1) `A new coffee shop can hold no more than 50 seats` 2) `The owner wants at least 20 of the seats to be stools and the remaining seats to be recliners` x= the number of stools y = number of recliners owner wants `no more than` 50 seats. for example, `a` is no more than `b` we write it as \[a \le b\] x+y should be less than or equal to 50 \[\rm x+y \le 50~~~~~(equation ~1)\] 2) at least 20 of the seats to be stools the inequality sign we use for `at least ` is ≥ \[x \ge 20~~~~(eq --> 2)\]

Nnesha:

https://www.desmos.com/ for graph